OpenStudy (anonymous):

-7x+5y=-68 2x+5y=-58 find the solution of the system of equations

OpenStudy (anonymous):

@Abhisar @Somy @summerschool

OpenStudy (anonymous):

subtract both sides of equation 1 from both sides of equation 2

OpenStudy (anonymous):

im stil kinda lost can u explain more this is new to me @Arnab09

OpenStudy (anonymous):

our goal is to bring the system down to one variable, either x or y. if we subtract left hand side(LHS) of the equation 1 from LHS of equation 2, what do we get? 2x+5y-(-7x+5y)=9x similarly, subtracting the RHS, we get -58-(-68)=10 so we are left with 9x=10 or x=10/9 put the value of x in either equation 1 or equation 2, you will find out y

OpenStudy (anonymous):

ohhhhh thnk very much

OpenStudy (anonymous):

welcome